tangents and normals
find the equations of all lines of slope zero and that are tangent to the curve y=1 / (x^2 – 2x + 3)

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

y = \frac{1}{x^2 - 2x + 3}
Slope of tangent: y’ = 0
y' = \frac{-(2x-2)}{(x^2 - 2x + 3)^2}
\frac{-(2x-2)}{(x^2 - 2x + 3)^2} = 0
\Rightarrow x = 1
Put in equation

y = \frac{1}{1^2 - 2.1 + 3}
y = \frac{1}{2}
this is the equation of tangent